If n=18, T(x-bar)-48, and s=4, find the margin of error at a 99% confidence level Give your answer to two decimal places.
Answer #1

Given, n-18, x=48, s=4 (203 Margin of error at a 991. confidence level is, ME = tn.id s x IS Where, 1-2=0.99 = 2 = 0.01 tn-1;Hi Dear,

I have given my 100% to solve your problem.

So, please help me out by just thumbs up.

Thank you, so much